OpenStudy (anonymous):

potassium-42 has a half life of 12,4 hours. How much of an 848 g sample of potassium-24 will be left after 62.0 hours?

OpenStudy (aaronq):

62 hours is how many half-lives?

OpenStudy (anonymous):

CAN YOU SHOW ME HOW YOU GOT THAT ANSWER?

OpenStudy (aaronq):

i didn't post any answers, i asked you how many half lives have passed in the question.

OpenStudy (anonymous):

62.0

OpenStudy (aaronq):

thats given in the question "potassium-42 has a half life of 12,4 hours. How much of an 848 g sample of potassium-24 will be left after \(\Huge\color{red}{62.0 \;hours}\)?"

OpenStudy (anonymous):

yes that is the given question

OpenStudy (aaronq):

so i repeat, 62 hours is how many half-lives?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

are you reading what I'm asking? how many half-lives have passed after 62 hours?

OpenStudy (anonymous):

You have 5 half-lives because 62/12.4= 5. Then you have to divide 848 g by 2 for 5 times... So let's do this out. First time: 848/2 = 424 Second: 424/2= 212 Third: 212/2= 106 Fourth: 53 Fifth: 26.5 Therefore, I believe your answer should be 26.5g of potassium-24